What is the p.m.f. (Probability Mass Function) for A1, where  A1 equals B1 minus B2 and we are given that Bi ∼ Ber(p).
Answer #1

PMF of Bi is given by,

P(Bi=0) = 1-p

P(Bi=1) = p

A1=B1-B2

A1 can take values -1,0 or 1

P(A1 = -1) = P(B1=0) * P(B2=1)

= (1-p)*p = p-p^2

P(A2 = 0) = P(B1=0) * P(B2=0) + P(B1=1) * P(B2=1)

= (1-p)*(1-p)+p*p = 1-2p+p^2+p^2 = 1-2p+2p^2

P(A1 = 1) = P(B1=1) * P(B2=0)

= p*(1-p) = p-p^2